OpenStudy (anonymous):

given the general identity tanx=sinx/cosx,which equation relating the acute angles,

OpenStudy (anonymous):

@AltCode

OpenStudy (anonymous):

can you help @Levity1

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

@phi please help?

OpenStudy (anonymous):

@Pawtpie @ParthKohli

OpenStudy (anonymous):

@Hero

hero (hero):

It's asking "Which equation" shows a relationship between angles A and C based on the identity which implies that you have a list of answers (you have not yet posted) to choose from.

OpenStudy (anonymous):

A. tanA=sinA/sinC B. cosA=tan(90-A)/sin(90-c) C.sinC=cosA/tanC D.cosA=tanC E. sinC=cos(90-c)/tanA

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

i am thinking it is A

hero (hero):

Can you explain using mathematical reasoning how you arrived at A as the answer?

OpenStudy (anonymous):

in all honesty it is a guess

OpenStudy (anonymous):

if its a right angle then it would be something like this

OpenStudy (anonymous):

\[\tan=\frac{ opposite }{ adjacent }\] \[Cos=\frac{ opposite }{ hypotenuse }\]

OpenStudy (anonymous):

right? \[\sin=\frac{ opposite }{ hypotenuse }\]

hero (hero):

That isn't precisely what explain it. But you have it correct. There's a general rule which says that if angles A and C are complementary (meaning if they add to 90 degrees) then sin(A) = cos(C) For example suppose A = 30° C = 60° Then sin(30°) = cos(60°) And sin(60°) = cos(30°) Therefore: \(\tan(30°) = \dfrac{\sin(30°)}{\sin(60°)} = \dfrac{\sin(30°)}{\cos(30°)}\) I hope that makes sense

OpenStudy (anonymous):

somewhat. but i still dont understand how to get the answer i need. The tan i have is \[tanA=\frac{ sinA }{ sinC}\]

hero (hero):

I just explained it. Read it ten more times.